OpenStudy (anonymous):

Solve. 6^2 + 8x = 14

OpenStudy (amistre64):

is it just luck that 6 and 14 subract to get 8?

OpenStudy (anonymous):

these are the choices for the answer x = 7/3 and x = 1 x = 7/3 and x = –1 x = 7/3– and x = –1 x =7/3 – and x = 1

OpenStudy (amistre64):

i assume thats spose to be 6x^2?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

6x^2 + 8x = 14 ; subtract 14 from each side -14 -14 ----------------- 6x^2 + 8x - 14 = 0 ; factor :) 6(14) = .... doesnt matter since 14-6 - 8 (x +14\6) (x -6\6) (x +7\3) (x -1) x = -7/3 and 1

OpenStudy (anonymous):

36 + 8x = 14 8x = -22 4x = -11 x (-11/4)